OpenStudy (anonymous):

Solve the system of equations by the substitution method. If the system does not have a single ordered pair as a solution, state whether the system is inconsistent or dependent. x + 3y = 32 -3x + 2y = 3 A. (4, 10) B. No solution; inconsistent C. (5, 9) D. (-5, 10)

OpenStudy (anonymous):

I don't think it's A

OpenStudy (kira_yamato):

x + 3y = 32 ----(1) -3x + 2y = 3 ---(2) From (1): x = 32 - 3y ---(3) Make use of substitution

OpenStudy (anonymous):

One sec

OpenStudy (kira_yamato):

I've solved it. Option A seems to be wrong...

OpenStudy (anonymous):

I'm solving it now give me a sec I have to use my tablet.

OpenStudy (anonymous):

Is it C?

OpenStudy (kira_yamato):

Yep

OpenStudy (anonymous):

Thank you sorry for taking so long

OpenStudy (kira_yamato):

No problems ^_^